OpenStudy (anonymous):

a sum of money doubles in 20 years on simple interest .it will get triple at the same rate in

OpenStudy (perl):

use the equation A = P(1 + rt)

OpenStudy (perl):

2P = P ( 1 + r*20) solve for r 2 = 1 + r*20 1 = r*20 r = 1/20 = .05 Now plug r = .05 into the formula A = P(1 + r*t) , and substitute r = .05, and t = 3 , A = 3P